# boxedasideAside

### v1

@@ A.2 The definition of $\overline{M}_{E,g}$ and $\overline{M}_{E,g;n}$
In this section, we define the Deligne-Mumford stack $\overline{M}_{E,g;n}$ as certain substack of $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E\times B\mathbb{Z}/2\mathbb{Z},1)$. In the rest of the paper, we require $g\geq 1,n\geq 0$, and $g+n\geq 2$ for $\overline{M}_{E,g;n}$, so that the object in the moduli problem has finite automorphism group.
@@
    We define $\widetilde{M}_{E,g;n}/\mathbb{S}$ as the open and closed substack of $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E\times B\mathbb{Z}/2\mathbb{Z},1)$ parameterizing genus 1 stable maps
@@
    We define $\overline{M}_{E,g;n}\coloneqq\widetilde{M}_{E,g;n}/S_{2g-2}\times S_{n}$, where $S_{2g-2}$ acts by permuting the first $22g-$ 2 marked points (the branching locus) and $S_{n}$ permutes the $n$ non-branching marked points.
@@
    When there is no unramified marked point (that is when $n=0$), we denote $\widetilde{M}_{E,g},\overline{M}_{E,g}$ and $M_{E,g}$ for the corresponding Deligne-Mumford stack.
@@
To unravel the definition, we describe the $k$-points of $\overline{M}_{E,g;n}$. Suppose $(\mathcal{C}\to{\rm Spec}(k),\Sigma_{{\rm br}},\Sigma_{{\rm ur}};f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ is a $k$-point of $\overline{M}_{E,g;n}$. To describe the map $f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z}$, we separately describe the content of the maps $\mathcal{C}\to E$ and $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$:
@@
    The map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$: By [ACV03, Theorem 4.3.2], this is equivalent to the data of an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ in the sense of [ACV03, Definition 4.3.1]. For the reader unfamiliar with admissible covers, an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ is essentially a $\mathbb{Z}/2\mathbb{Z}$-cover where we only allow branching above the marked points and nodes. Because we require the stabilizers of $\mathcal{C}$ at $\Sigma_{{\rm br}}$ to be nontrivial and the map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$ to be representable, the $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ must be branched over $\Sigma_{{\rm br}}$. Similarly, the stabilizers of $\mathcal{C}$ at $\Sigma_{{\rm ur}}$ are trivial, so it is unbranched over $\Sigma_{{\rm ur}}$.
    If $g\geq 2$, then the fact that there are $22g-$ 2 branch points ensures that the genus of $D$ is $g$ by the Riemann-Hurwitz formula. We discuss the $g=1$ case in Remark A.7.
In summary, a $k$-point $(\mathcal{C}\to{\rm Spec}k,\Sigma_{{\rm br}},\Sigma_{{\rm ur}},f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ of $\overline{M}_{E,g;n}$ is the same data as a nodal curve $C$ with one component identified with $E$ and all other components genus 0 bubbles, along with an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $\pi:D\to C$ branched above $\Sigma_{{\rm br}}$ and possibly the nodes. The $k$-point lies in $M_{E,g;n}$ if $C=E$. Note that because $g+n\geq 2$, there must be at least one marked point, and $(C,\Sigma_{{\rm br}}\cup\Sigma_{{\rm ur}})$ must actually be stable. This implies that $(D,\Sigma_{{\rm br}}^{\prime}\cup\Sigma_{{\rm ur}}^{\prime})$ is also stable, where $\Sigma_{{\rm br}}^{\prime}$ is the divisor of ramification points over the marked points $\Sigma_{{\rm br}}$ (i.e. $\Sigma_{{\rm br}}^{\prime}=\pi^{-1}(\Sigma_{{\rm br}})_{\mathrm{red}}$) and $\Sigma_{{\rm ur}}^{\prime}=\pi^{-1}(\Sigma_{{\rm ur}})_{\mathrm{red}}$.
@@
We take a moment to describe the points of $\overline{M}_{E,1}$, which correspond to étale double covers of $E$. There are $4$tale double covers. Suppose $E[2](k)=\{O,P_{1},P_{2},P_{3}\}$, then the étale covers of $E$ are of the following two types:
@@
A variation of the above construction where we consider stable maps to $C\times BG$ can be used to construct spaces of admissible $G$-covers of a fixed base curve $C$ for arbitrary $G$ and $C$, over a base scheme $\mathcal{S}$ such that #$\#G$ is coprime to the characteristic of $\mathcal{O}_{S}$.
@@ A.3 Deformation theory of $\overline{M}_{E,g;n}$
Let $(C\to{\rm Spec}k,\Sigma_{{\rm br}}\cup\Sigma_{ur},f:C\to E)$ be a stable map corresponding to an object of $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E,1)(k)$. We want to show that this object has no nontrivial infinitesimal automorphisms. If $C$ has more than one component, then it is stable, and $C$ automatically has no nontrivial infinitesimal automorphisms. Otherwise, $C\cong E$, and only the trivial infinitesimal automorphism can preserve the isomorphism $f$. Thus, $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E,1)$ has no infinitesimal automorphisms and is Deligne-Mumford. ∎
We use the deformation-obstruction theory for stable maps in [AV02, §5.3] to prove that $\overline{M}_{E,g;n}$ is smooth and compute its dimension. Throughout the following proof, if $f:X\to Y$ is a map, we will use $\mathbb{L}_{X/Y}$ to denote the relative cotangent complex. If $Y={\rm Spec}k$, we will just write $\mathbb{L}_{X}$.
@@
$\overline{M}_{E,g;n}$ is smooth of dimension $2g-2+n$.
@@
Obstructions: First, we show that $\overline{M}_{E,g;n}$ is smooth by showing the obstruction groups vanish. Let $A^{\prime}\to A$ be a small extension of local Artinian $k$-algebras with $I\coloneqq\mathrm{ker}(A^{\prime}\to A)$, and let $(\mathcal{C}\to{\rm Spec}A,\Sigma_{{\rm br}}\cup\Sigma_{{\rm ur}},f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be an object of $\overline{M}_{E,g;n}(A)$. Let $\mathcal{C}_{0}\coloneqq\mathcal{C}\times_{{\rm Spec}A}{\rm Spec}k$. By [AV02, Lemma 5.3.3], the obstruction to lifting this object to $A^{\prime}$ lies in ${\rm Ext}_{\mathcal{O}_{C}}^{2}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2\mathbb{Z}},\mathcal{O}_{\mathcal{C}_{0}})\otimes_{k}I\cong{\rm Ext}_{\mathcal{O}_{C_{0}}}^{2}(\mathbb{L}_{\mathcal{C}_{0}/E\times B\mathbb{Z}/2\mathbb{Z}},\mathcal{O}_{\mathcal{C}_{0}})\otimes_{k}I$. For brevity, we will write ${\rm Ext}$ to mean ${\rm Ext}_{\mathcal{O}_{\mathcal{C}_{0}}}$ for the rest of this proof.
@@
    is injective. Let $s\in H^{0}(\mathcal{C}_{0},f^{*}\Omega_{E}\otimes\omega_{\mathcal{C}_{0}})$ be a nonzero section. Because the dualizing sheaf is étale local [Nir08, Lemma 2.12], we can identify $\omega_{\mathcal{C}_{0}}$ with the sheaf of 1-forms with logarithmic singularities whose residues at both components of each node add to 0. In particular, any section of $f^{*}\Omega_{E}\otimes\omega_{\mathcal{C}_{0}}$ whose restriction to some component is nonzero must be nonzero on each component. This means that $s$ must be nonzero when restricted to the component of $\mathcal{C}_{0}$ whose coarse space is identified with $E$. On this component, the map $f^{*}\Omega_{E}\to\Omega_{\mathcal{C}_{0}}$ is generically an isomorphism and hence injective, which means that the image $\psi(s)$ is nonzero. We conclude that $\psi$ is injective and that the given map is surjective.
@@
Deformations: We compute the dimension of $\overline{M}_{E,g;n}$ by computing the dimension of tangent space at every point. Let $(\mathcal{C}\to{\rm Spec}k,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be a object of $\overline{M}_{E,g;n}(k)$. Then by [AV02, Lemma 5.3.2], the tangent space $T$ of $\overline{M}_{E,g;n}$ at the point representing this object sits in the following exact sequence of vector spaces:
$0\to\text{Hom}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})\to H^{0}(\mathcal{C},\mathcal{N}_{\Sigma})\to T\to{\rm Ext}^{1}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})\to 0$
The exact sequence stated in [AV02, Lemma 5.3.2] actually only gives us the above sequence from the second term $\text{Hom}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})$ onwards. However, as in [ACV03, §3.0.4], we extend this exact sequence so that the first term is the infinitesimal automorphism group of the data $(\mathcal{C}\to{\rm Spec}k,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$. By stability, this infinitesimal automorphism group vanishes, so we get the above exact sequence starting with 0.
First notice that $\mathcal{N}_{\Sigma}=\oplus_{i}\mathcal{N}_{\Sigma_{i}}$, where $\mathcal{N}_{\Sigma_{i}}$ is the normal sheaf at the $i^{th}$ marked point. By [ACV03, 3.0.4], at each $\sigma_{i}$ where the marked is a branching point, then $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma_{i}})=0$. At each $\tau_{j}$ where the mark is not a branching point, then $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma_{j}})\cong k$. Therefore, $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma})\cong k^{n}$.
@@
$f^{*}\Omega_{E}\to\Omega_{\mathcal{C}}\to\mathbb{L}_{\mathcal{C}/E}$
That gives:
@@
Since as above ${\rm Ext}^{2}(\mathbb{L}_{\mathcal{C}/E},\mathcal{O}_{C})=0$. So the dimension of $T$ is
@@
by the exact sequence (7). Since $\Omega_{E}$ is trivial, $\text{Hom}(f^{*}\Omega_{E},\mathcal{O}_{\mathcal{C}})\cong H^{0}(\mathcal{C},\mathcal{O}_{\mathcal{C}})=H^{0}(C,\mathcal{O}_{C})\cong k$, and ${\rm Ext}^{1}(f^{*}\Omega_{E},\mathcal{O}_{\mathcal{C}})\cong H^{1}(\mathcal{C},\mathcal{O}_{\mathcal{C}})=H^{1}(C,\mathcal{O}_{C})\cong k$ since $C$ has genus $1$.
@@
Since all of our nodes are balanced, we have that $\pi_{*}(\Omega_{\mathcal{C}}\otimes\omega_{\mathcal{C}})=\Omega_{C}\otimes\omega_{C}(\Sigma_{\rm br})$ as in [ACV03, §3.0.4], where $\Sigma_{\rm br}$ is summing over all the smooth branched markings $\sigma_{i}$. Therefore,
@@
    The construction of $X$: Let $\mathcal{L}$ be the universal bundle on $E\times\textup{Pic}^{g-1}(E)$, and let $\pi:E\times\textup{Pic}^{g-1}(E)\to\textup{Pic}^{g-1}(E)$ be the projection. Let $\mathcal{V}\coloneqq\pi_{*}\mathcal{L}^{\otimes 2}$. Because $02g-2>$ 0 and $H^{1}(E,\mathcal{M})\cong 0$ for any positive degree line bundle $\mathcal{M}$, by Serre duality, cohomology and base change tells us that $\mathcal{V}$ is a locally constant sheaf on $\textup{Pic}^{g-1}(E)$.
    By abuse of notation, let $\mathcal{V}$ also denote the space of global sections of $\mathcal{V}$ (as we will do throughout this proof), we let $X\subset\mathcal{V}$ be the open subset consisting of sections whose vanishing loci are reduced. In other words, $X$ consists of pairs $(\mathcal{M},r)$, where $\mathcal{M}$ is a line bundle on $E$ of degree $g-1$ and the zeroes of $r\in H^{0}(E,\mathcal{M}^{\otimes 2})$ have multiplicity 1. Because $\textup{Pic}^{g-1}(E)$ is irreducible, $X$ is irreducible.

### v2

@@ A.2 The definition of $\overline{M}_{E,g}$ and $\overline{M}_{E,g;n}$
In this section, we define the Deligne-Mumford stack $\overline{M}_{E,g;n}$ as certain substack of the stack $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E\times B\mathbb{Z}/2\mathbb{Z},1)$. In the rest of the paper, we require $g\geq 1,n\geq 0$, and $g+n\geq 2$ for $\overline{M}_{E,g;n}$, so that the object in the moduli problem has finite automorphism group.
@@
    We define $\widetilde{M}_{E,g;n}/\mathbb{S}$ as the open and closed substack of $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E\times B\mathbb{Z}/2\mathbb{Z},1)$ parametrizing genus 1 stable maps
@@
    We define $\overline{M}_{E,g;n}\coloneqq[\widetilde{M}_{E,g;n}/S_{2g-2}\times S_{n}]$, where $S_{2g-2}$ acts by permuting the first 2g22g-2 marked points (the branching locus) and $S_{n}$ permutes the $n$ unbranched marked points.
@@
    When there is no unbranched marked point (that is when $n=0$), we denote $\widetilde{M}_{E,g},\overline{M}_{E,g}$ and $M_{E,g}$ for the corresponding Deligne-Mumford stack.
@@
To unravel the definition of $\overline{M}_{E,g;n}$, we describe its $k$-points. To this end, suppose $(\mathcal{C}\to{\rm Spec}(k),\Sigma_{{\rm br}},\Sigma_{{\rm ub}};f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ is a $k$-point of $\overline{M}_{E,g;n}$. To describe the map $f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z}$, we separately describe the content of the maps $\mathcal{C}\to E$ and $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$:
@@
    The map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$: By [ACV03, Theorem 4.3.2], this is equivalent to the data of an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ in the sense of [ACV03, Definition 4.3.1]. For the reader unfamiliar with admissible covers, an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ is essentially a $\mathbb{Z}/2\mathbb{Z}$-cover where we only allow branching above the marked points and nodes. Because we require the stabilizers of $\mathcal{C}$ at $\Sigma_{{\rm br}}$ to be nontrivial and the map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$ to be representable, the $\mathbb{Z}/2\mathbb{Z}$-cover $D\to C$ must be branched over $\Sigma_{{\rm br}}$. Similarly, the stabilizers of $\mathcal{C}$ at $\Sigma_{{\rm ub}}$ are trivial, so it is unbranched over $\Sigma_{{\rm ub}}$.
    If $g\geq 2$, then the fact that there are 2g22g-2 branch points ensures that the genus of DD is $g$ by the Riemann-Hurwitz formula. We discuss the $g=1$ case in Remark A.8.
In summary, a $k$-point $(\mathcal{C}\to{\rm Spec}(k),\Sigma_{{\rm br}},\Sigma_{{\rm ub}},f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ of $\overline{M}_{E,g;n}$ is the same data as a nodal curve $C$ with one component identified with $E$ and all other components genus 0 bubbles, along with an admissible $\mathbb{Z}/2\mathbb{Z}$-cover $\pi:D\to C$ branched above $\Sigma_{{\rm br}}$ and possibly the nodes. The $k$-point lies in $M_{E,g;n}$ if $C=E$. Note that because $g+n\geq 2$, there must be at least one marked point, and $(C,\Sigma_{{\rm br}}\cup\Sigma_{{\rm ub}})$ must actually be stable. This implies that $(D,\Sigma_{{\rm br}}^{\prime}\cup\Sigma_{{\rm ub}}^{\prime})$ is also stable, where $\Sigma_{{\rm br}}^{\prime}$ is the divisor of ramification points over the marked points $\Sigma_{{\rm br}}$ (i.e. $\Sigma_{{\rm br}}^{\prime}=\pi^{-1}(\Sigma_{{\rm br}})_{\mathrm{red}}$) and $\Sigma_{{\rm ub}}^{\prime}=\pi^{-1}(\Sigma_{{\rm ub}})_{\mathrm{red}}$.
@@
Instead of constructing $\overline{\mathcal{B}}_{E,g}$ as the image of the forgetful map $\overline{M}_{E,g}\to\overline{\mathcal{M}}_{g}$, one could alternatively use the stack $\overline{\mathcal{H}}_{g,G,\xi}$ of admissible double covers of genus 1 curves, as defined in [SvZ20, §1.3], with Galois group $G=\mathbb{Z}/2\mathbb{Z}$ and ramification $\xi=(1,\ldots,1)$. In Section 3.3, we denote $\overline{\mathcal{B}}_{g}^{\rm adm}=\overline{\mathcal{H}}_{g,\mathbb{Z}/2\mathbb{Z},\xi}$. Recall from Section 3.3 that $\overline{\mathcal{B}}_{g}^{\rm adm}$ admits a map $\Psi:\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{g}$, extracting the source curve, and a map $\Phi:\overline{\mathcal{B}}_{g}^{\rm adm}\to\overline{\mathcal{M}}_{1,1}$, extracting the target curve with one marked point. After fixing a point $E]=(E,O_{E})\in\overline{\mathcal{M}}_{1,1}$, we may alternatively define $\overline{\mathcal{B}}_{E,g}=\Psi(\Phi^{-1}([E]))$. The fact that this definition is equivalent to our definition follows from [SvZ20, Theorem 3.7] and Proposition 2.9. The reason we instead construct $\overline{M}_{E,g}$ from scratch and realize $\overline{\mathcal{B}}_{E,g}$ as the image of $\overline{M}_{E,g}$ under the forgetful map is that our construction allows us to prove that $\overline{\mathcal{B}}_{E,g}$ is irreducible.
Remark A.8.
We take a moment to describe the points of $\overline{M}_{E,1}$, which correspond to étale double covers of $E$. There are $4$ étale double covers. Suppose $E[2](k)=\{O,P_{1},P_{2},P_{3}\}$, then the étale covers of $E$ are of the following two types:
@@
A variation of the above construction where we consider stable maps to $C\times BG$ can be used to construct spaces of admissible $G$-covers of a fixed base curve $C$ for arbitrary $G$ and $C$, over a base scheme $\mathcal{S}$ such that #G\#G is coprime to the characteristic of $\mathcal{O}_{S}$.
@@ A.3 Deformation theory of $\overline{M}_{E,g;n}$
Let $(C\to{\rm Spec}(k),\Sigma_{{\rm br}}\cup\Sigma_{{\rm ub}},f:C\to E)$ be a stable map corresponding to an object of $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E,1)(k)$. We want to show that this object has no nontrivial infinitesimal automorphisms. If $C$ has more than one component, then it is stable, and $C$ automatically has no nontrivial infinitesimal automorphisms. Otherwise, $C\cong E$, and only the trivial infinitesimal automorphism can preserve the isomorphism $f$. Thus, $\mathcal{K}_{1,2g-2+n}^{\text{bal}}(E,1)$ has no infinitesimal automorphisms and is Deligne-Mumford. ∎
We use the deformation-obstruction theory for stable maps in [AV02, §5.3] to prove that $\overline{M}_{E,g;n}$ is smooth and compute its dimension. Throughout the following proof, if $f:X\to Y$ is a map, we will use $\mathbb{L}_{X/Y}$ to denote the relative cotangent complex. If $Y={\rm Spec}(k)$, we will just write $\mathbb{L}_{X}$.
@@
$\overline{M}_{E,g;n}$ is smooth of dimension 2g2+n2g-2+n.
@@
Obstructions: First, we show that $\overline{M}_{E,g;n}$ is smooth by showing the obstruction groups vanish. Let $A^{\prime}\to A$ be a small extension of local Artinian $k$-algebras with $I\coloneqq\mathrm{ker}(A^{\prime}\to A)$, and let $(\mathcal{C}\to{\rm Spec}(A),\Sigma_{{\rm br}}\cup\Sigma_{{\rm ub}},f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be an object of $\overline{M}_{E,g;n}(A)$. Moreover, define $\mathcal{C}_{0}\coloneqq\mathcal{C}\times_{{\rm Spec}(A)}{\rm Spec}(k)$. By [AV02, Lemma 5.3.3], the obstruction to lifting this object to $A^{\prime}$ lies in ${\rm Ext}_{\mathcal{O}_{C}}^{2}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2\mathbb{Z}},\mathcal{O}_{\mathcal{C}_{0}})\otimes_{k}I\cong{\rm Ext}_{\mathcal{O}_{C_{0}}}^{2}(\mathbb{L}_{\mathcal{C}_{0}/E\times B\mathbb{Z}/2\mathbb{Z}},\mathcal{O}_{\mathcal{C}_{0}})\otimes_{k}I$. For brevity, we will write $\rm Ext$} to mean ${\rm Ext}_{\mathcal{O}_{\mathcal{C}_{0}}}$ for the rest of this proof.
@@
    is injective. Let $s\in H^{0}(\mathcal{C}_{0},f^{*}\Omega_{E}\otimes\omega_{\mathcal{C}_{0}})$ be a nonzero section. Because the dualizing sheaf is étale local [Nir08, Lemma 2.12], we can identify $\omega_{\mathcal{C}_{0}}$ with the sheaf of 1-forms with logarithmic singularities whose residues at both components of each node add to 0. In particular, any section of $f^{*}\Omega_{E}\otimes\omega_{\mathcal{C}_{0}}$ whose restriction to some component is nonzero must be nonzero on each component. This means that $s$ must be nonzero when restricted to the component of $\mathcal{C}_{0}$ whose coarse space is identified with $E$. On this component, the map $f^{*}\Omega_{E}\to\Omega_{\mathcal{C}_{0}}$ is generically an isomorphism and hence injective, which means that the image $\psi(s)$ is nonzero. We conclude that $\psi$ is injective and that ${\rm Ext}^{1}(\Omega_{\mathcal{C}_{0}},\mathcal{O}_{\mathcal{C}_{0}})\to{\rm Ext}^{1}(f^{*}\Omega_{E},\mathcal{O}_{\mathcal{C}_{0}})$ is surjective.
@@
Deformations: We compute the dimension of $\overline{M}_{E,g;n}$ by computing the dimension of tangent space at every point. Let $(\mathcal{C}\to{\rm Spec}(k),\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be a object of $\overline{M}_{E,g;n}(k)$. Then by [AV02, Lemma 5.3.2], the tangent space $T$ of $\overline{M}_{E,g;n}$ at the point representing this object sits in the following exact sequence of vector spaces:
$0\to\text{Hom}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})\to H^{0}(\mathcal{C},\mathcal{N}_{\Sigma})\to T\to{\rm Ext}^{1}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})\to 0.$
The exact sequence stated in [AV02, Lemma 5.3.2] actually only gives us the above sequence from the second term $\text{Hom}(\mathbb{L}_{\mathcal{C}/E\times B\mathbb{Z}/2},\mathcal{O}_{\mathcal{C}})$ onwards. However, as in [ACV03, §3.0.4], we extend this exact sequence so that the first term is the infinitesimal automorphism group of the data $(\mathcal{C}\to{\rm Spec}(k),\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$. By stability, this infinitesimal automorphism group vanishes, so we get the above exact sequence starting with 0.
First notice that $\mathcal{N}_{\Sigma}=\oplus_{i}\mathcal{N}_{\Sigma_{i}}$, where $\mathcal{N}_{\Sigma_{i}}$ is the normal sheaf at the $i^{th}$ marked point. By [ACV03, 3.0.4], at each $\sigma_{i}$ where the marked is a branching point, then $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma_{i}})=0$. At each $\tau_{j}$ where the marked point is not a branching point, we have $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma_{j}})\cong k$. Therefore, we obtain $H^{0}(\mathcal{C},\mathcal{N}_{\Sigma})\cong k^{n}$.
@@
$f^{*}\Omega_{E}\to\Omega_{\mathcal{C}}\to\mathbb{L}_{\mathcal{C}/E}.$
This gives the long exact sequence
@@
As above, we use ${\rm Ext}^{2}(\mathbb{L}_{\mathcal{C}/E},\mathcal{O}_{C})=0$. Hence, the dimension of $T$ is
@@
by the exact sequence (9). Since $\Omega_{E}$ is trivial and $C$ has genus $1$, we observe
$\text{Hom}(f^{*}\Omega_{E},\mathcal{O}_{\mathcal{C}})$ $\cong H^{0}(\mathcal{C},\mathcal{O}_{\mathcal{C}})=H^{0}(C,\mathcal{O}_{C})\cong k$
${\rm Ext}^{1}(f^{*}\Omega_{E},\mathcal{O}_{\mathcal{C}})$ $\cong H^{1}(\mathcal{C},\mathcal{O}_{\mathcal{C}})=H^{1}(C,\mathcal{O}_{C})\cong k.$
@@
Since all of our nodes are balanced, we have $\pi_{*}(\Omega_{\mathcal{C}}\otimes\omega_{\mathcal{C}})=\Omega_{C}\otimes\omega_{C}(\Sigma_{\rm br})$ as in [ACV03, §3.0.4], where $\Sigma_{\rm br}$ is summing over all the smooth branched markings $\sigma_{i}$. Therefore,
@@
    The construction of $X$: Let $\mathcal{L}$ be the universal bundle on $E\times\textup{Pic}^{g-1}(E)$, and consider the projection $\pi:E\times\textup{Pic}^{g-1}(E)\to\textup{Pic}^{g-1}(E)$. Let $\mathcal{V}\coloneqq\pi_{*}\mathcal{L}^{\otimes 2}$. Because 2g2>02g-2>0 and $H^{1}(E,\mathcal{M})\cong 0$ for any positive degree line bundle $\mathcal{M}$, by Serre duality, cohomology and base change tells us that $\mathcal{V}$ is a locally constant sheaf on $\textup{Pic}^{g-1}(E)$.
    By abuse of notation, let $\mathcal{V}$ also denote the space of global sections of $\mathcal{V}$ (as we will do throughout this proof). We let $X\subset\mathcal{V}$ be the open subset consisting of sections whose vanishing loci are reduced. In other words, $X$ consists of pairs $(\mathcal{M},r)$, where $\mathcal{M}$ is a line bundle on $E$ of degree g1g-1 and the zeroes of $r\in H^{0}(E,\mathcal{M}^{\otimes 2})$ have multiplicity 1. Because $\textup{Pic}^{g-1}(E)$ is irreducible, $X$ is irreducible.